OpenStudy (anonymous):

Write an equation for a line in slope-intercept form that passes through the points (3, 4) and (1, -2).

jimthompson5910 (jim_thompson5910):

Equation is \[\large y=3x-5\] See attached image for work

OpenStudy (aroub):

-2-4/1-3 -6/-2 3 y=mx+b , m=3 , (3,4) 4=3(3)+b 4=9+b -9+4=b -5=b y=3x-5

OpenStudy (lalaly):

slope =m= y2-y1/x2-x1 = -2-4/1-3 = -6/-2=3 equation y-4=3(x-3) y=3x-5